### v1

@@ VI-F Negative Transfer and Overfitting
VII New Perspectives and Current Trends
The results of previous research indicate that, despite having achieved high levels of accuracy, there is still room for improvement in the determination of KC. To that end, scientists propose using a less common set of algorithms that can aid in accurately identifying these cells, especially in their early stages. Enhancing the capability of models to accurately determine the presence of KC at various stages while preserving accuracy and improving unbiasedness is crucial. The combination of feature selection methods and DL models may prove helpful in this regard. A larger database is required, and the algorithm must be refined and optimized for other classification tasks related to KC detection. The data collected from various medical institutions and even available databases can be used to evaluate these strategies. There are several potential future directions for the use of TL in the context of KC diagnosis:
VII-A Multimodal Learning
Multimodal learning represents a strategy that integrates information from diverse sources or data types to enhance learning and decision-making processes (Fig.6). Within cancer diagnosis, this approach entails the amalgamation of data from varied origins, including medical imaging (such as MRI and CT scans), genomic data, and clinical records (like patient history and laboratory results). Merging multimodal learning with TL for cancer diagnosis involves training a comprehensive model on a dataset encompassing medical images, genomic information, and clinical data across multiple cancer varieties. Subsequently, TL techniques are employed to refine the model for diagnosing a particular cancer type. For instance, a DL model could initially be educated on a broad array of medical images from different cancers, with TL utilized to tailor the model for specific cancer types. Furthermore, the model might also undergo training with genomic and clinical data, integrating these disparate data types through multimodal learning methods to enhance the model’s precision and robustness. In essence, the application of multimodal learning in conjunction with TL for cancer diagnosis promises to elevate diagnostic accuracy, diminish instances of false positives and negatives, and facilitate earlier, more precise cancer detection.
Figure 6: Multimodal learning for KC
VII-B Federated Learning
Federated learning (FL) [175] is gaining momentum in the field of KC diagnosis, prediction, and segmentation, demonstrating substantial potential in enhancing accuracy and preserving data privacy across various healthcare environments. This collaborative approach enables the utilization of data from multiple institutions without compromising patient confidentiality, addressing both the variability and volume of medical data essential for effective machine-learning models. In the study by Vekaria et al. [176], the integration of FL with TL is explored as a novel approach to enhance the detection of kidney abnormalities from medical imaging. This research demonstrates how FL can facilitate collaborative learning across different healthcare institutions without necessitating direct data sharing. By doing so, it not only ensures data privacy but also leverages diverse datasets to improve diagnostic accuracy and model robustness. Gao et al. [177] highlight the effectiveness of a federated learning-driven collaborative diagnostic system for metastatic breast cancer. The ability of this system to improve generalizability and diagnostic accuracy without compromising patient privacy offers a compelling model for KC diagnostics, particularly in settings with diverse patient populations and imaging protocols. Similarly, the work by Subashchandrabose et al. [178] introduces an ensemble-based FL framework for the classification of lung cancer, which can be analogously applied to KC. This approach integrates multiple machine learning models trained on disparate datasets to enhance both accuracy and generalizability. The success of this method underscores the potential of FL in overcoming the limitations posed by single-institution datasets, suggesting a scalable model for KC diagnosis that respects patient privacy.
Further extending the application of FL, Ma et al. [179] develop an assisted diagnosis model for cancer recurrence using federated DL. This model not only predicts the recurrence of cancer but also identifies potential recurrence locations, offering a significant tool for post-operative care and monitoring in KC patients. By incorporating various physical examination indicators and leveraging a federated architecture, this approach enhances prediction accuracy while safeguarding sensitive patient data. Moving forward, Yaqoob et al. [180] focus on the implementation challenges of traditional machine learning in skin cancer but their insights are equally applicable to KC. They highlight the advantages of FL in building privacy-aware healthcare systems that comply with stringent data privacy regulations. This review provides a solid foundation for understanding how FL can be adapted for KC diagnosis and prediction, ensuring data privacy without compromising on diagnostic capabilities.
Tan et al. [181] and Li et al. [182] explore advanced FL frameworks that incorporate TL and data augmentation techniques like SMOTE to enhance the detection and classification capabilities in medical imaging. These studies illustrate the effectiveness of FL in handling imbalanced datasets, which is a common issue in KC diagnosis due to the variability of lesion appearances and stages.
A study by Kumbhare et al. [183] provides an insightful framework utilizing FL for breast cancer detection, which can be adapted for KC. This study outlines an innovative AI training method that allows hospitals to collectively harness the power of vast datasets from multiple sources to improve diagnosis accuracy. The approach not only respects patient privacy but also reduces processing times, indicating a promising avenue for KC applications where early detection and accurate segmentation are crucial. Almufareh et al. [184] demonstrate the use of FL in developing a highly accurate breast cancer detection model. By utilizing DNNs and data augmentation techniques within a federated framework, this study shows significant advancements in model performance. The methodology promises applicability to KC, where complex imaging features require sophisticated model tuning and privacy-preserving collaborative efforts. Jimenez et al. [185] explore the memory-aware curriculum learning in a federated setting to improve the classification of breast cancer across different clinical sites. This approach, focusing on optimizing the order and frequency of local updates, could be particularly beneficial for KC diagnosis, where data from multiple institutions vary significantly.
Lastly, Heidari et al. [186] describe the use of blockchain-based FL for detecting malignant lung nodules, showcasing the method’s robustness in preserving data integrity and anonymity across collaborating entities. This model’s success in maintaining privacy while achieving high accuracy underscores its potential adaptability to KC, where similar challenges in data variability and privacy concerns are prevalent.
VII-C Self-Paced Learning (SPL)
Self-paced learning (SPL) has emerged as a potent methodology for enhancing medical diagnostic systems, particularly in the challenging arena of kidney cancer diagnosis. This learning strategy prioritizes learning samples based on their complexity, adapting the training process to focus progressively on the more difficult or informative cases as the model’s capability improves. This adaptive learning approach can significantly benefit the diagnosis of kidney cancer, where data variability and imbalance often pose significant challenges. One significant application of self-paced learning in the medical field is addressed by Wang et al. (2020), who developed a self-paced learning sampling method specifically designed for imbalanced cancer classification. This method helps to reduce noise from imbalanced samples, improving the pre-diagnosis accuracy of cancer data, which is crucial for conditions like kidney cancer where early detection can significantly influence treatment outcomes [187]. Furthermore, the integration of SPL with federated learning, as discussed by Wang and Zhou, creates a privacy-preserving disease diagnosis framework [188]. This combination is particularly pertinent in decentralized healthcare systems where data privacy is crucial but collaboration across institutions is needed to enhance diagnostic accuracy. The authors in [189] also introduced a Random Forest with Self-Paced Bootstrap Learning for lung cancer prognosis, which can be adapted for kidney cancer. This framework utilizes self-paced learning to enhance the accuracy of cancer diagnosis and classification, proving that SPL can effectively handle complex datasets and improve predictive performance.
The method developed by Huang and Liang, which integrates self-paced learning for gene expression analysis, could be similarly beneficial for kidney cancer diagnosis [190] Ref4. By allowing for the gradual inclusion of challenging samples into the training set, this approach aids in identifying new, reliable biomarkers for kidney cancer, facilitating targeted therapy and personalized medicine. Shen et al. (2019) demonstrated the utility of combining deep active learning with self-paced learning for breast mass detection from mammograms [191]. This methodology can be translated into kidney cancer diagnosis, where imaging plays a crucial role, helping radiologists to detect and categorize renal masses more effectively.
In practical applications, the deep active self-paced learning model by Wang et al. [192] for accurate pulmonary nodule segmentation exemplifies how SPL can refine learning algorithms to better segment and diagnose lung nodules, a methodology that can similarly benefit kidney tumor segmentation. This approach emphasizes the active learner’s role in selecting challenging or uncertain samples, thereby optimizing the learning process and outcomes in medical image analysis. The overarching potential of SPL in medical imaging is further supported by Li et al. [193], who developed a self-paced convolutional neural network for aiding detection in medical imaging analysis. This approach could be instrumental in refining training samples for kidney cancer imaging, improving the precision of computer-aided detection systems.
Wang et al. [194] explore deep active self-paced learning for biomedical imaging, particularly focusing on pulmonary nodule segmentation in lung cancer. This method, which adjusts training focus from simple to complex samples, could similarly enhance kidney cancer diagnosis by improving model accuracy on atypical tumor presentations. Asare et al. [195] discuss a self-training scheme that incorporates self-paced learning for skin lesion classification. This approach gradually introduces harder cases into the training set, which could help develop robust diagnostic models for kidney cancer, adapting to the disease’s varied manifestations. Lastly, [196] focuses on balancing the training process between easy and challenging cases in skin disease recognition, which could be applied to kidney cancer to enhance detection accuracy and reduce biases towards common presentations.
VII-D Blockchain Technology
Blockchain technology represents a decentralized ledger that chronicles an ever-expanding list of entries known as blocks, safeguarded through cryptographic means [197]. Within the realm of cancer diagnosis, TL can be employed to educate a DL model using a vast collection of medical imagery, including CT scans, MRIs, or X-rays, for the identification of malignant tumors. After initial training, this model can undergo refinement on a narrower dataset of cancer patients to enhance its diagnostic accuracy for a particular cancer variant. A significant hurdle in deploying DL models for cancer diagnostics is the requisite for accessing and analyzing substantial volumes of sensitive patient information, all while safeguarding patient confidentiality and data integrity. Blockchain technology emerges as a pivotal solution in this context. It enables the secure, immutable storage of medical data, thus ensuring transparency and accountability for patients, medical practitioners, and researchers alike. For example, it allows patients to exercise control over their medical records, offering permissions to specific researchers or doctors, while enabling physicians to verify the data’s authenticity and precision. Moreover, blockchain facilitates the exchange and cooperative use of medical data among various healthcare entities, potentially leading to more precise diagnoses and enhanced patient care outcomes [37]. Fig.7 show the blockchain learning for KC.
Figure 7: Blockchain learning for KC
VII-E Explainable AI
The potential of explainable AI (XAI) in the diagnosis and segmentation of KC presents a promising frontier in medical imaging and oncology [198]. The challenge of interpreting complex diagnostic images, such as kidney-ureter-bladder (KUB) X-ray images, is well-documented. Ahmed et al. [199] address this by applying a TL model with a pre-trained VGG16 architecture, enhanced with XAI techniques like Layer-Wise Relevance Propagation (LRP), achieving a remarkable 97.41% accuracy in classifying KUB X-ray images as kidney stones or normal cases. This incorporation of XAI not only boosts the model’s performance but also its transparency, making the decision-making process more understandable to medical professionals. Similarly, Wani et al. [200] introduce ”DeepXplainer”, a hybrid DL model for lung cancer detection that employs SHapley Additive exPlanations (SHAP) for explicability. Although focused on lung cancer, the principles outlined could be adapted for KC diagnosis, highlighting the versatility of XAI in enhancing trust and comprehension of AI predictions in healthcare.
In kidney disease management, Moreno et al. [201] developed an explainable AI model for early detection of Chronic Kidney Disease (CKD), demonstrating how XAI can elucidate the contributions of clinical features like hemoglobin and hypertension in diagnosis. This approach can be similarly applied to KC, where understanding the influence of specific radiological features on diagnostic outcomes can significantly enhance early detection and treatment strategies. Arumugham et al. [202] and Azad et al. [203] both highlight the use of DL models combined with XAI techniques such as LIME and GRAD-CAM to improve the interpretability of predictions in kidney disease and cancer detection from various imaging modalities. These methods not only provide a heatmap or visual explanation of the AI’s focus during analysis but also reassure clinicians about the AI’s reliability, fostering a greater trust in AI-driven diagnostics. Additionally, Takahashi et al. [204] focus on the use of CNNs for detecting KC from CT images, where they employ Grad-CAM for visual explanations. This technique can be particularly useful in KC segmentation, as it highlights areas of interest that lead to a diagnosis, aiding radiologists in understanding and verifying the AI-generated outputs.
In the context of KC prediction, Lazebnik et al. [205] predict acute kidney injury (AKI) post-partial nephrectomy using machine learning algorithms with explainability components, achieving 75% accuracy. This approach underscores the critical role of preoperative parameters in predicting postoperative complications, with XAI offering a pathway to understand and improve these predictive models. In the same direction, Han et al. (2021) tackle the prediction of renal cell carcinoma (RCC) using deep convolutional TL and SHAP for model interpretability, achieving significant accuracies and providing valuable insights into model decisions [206]. This approach not only facilitates the understanding of how certain features affect the model’s prediction but also improves the clinical utility of the models by providing insights that are actionable in clinical practice.
Lastly, the research by Javaheri et al. [207] on differentiating between renal oncocytoma and chromophobe renal cell carcinoma using explainable AI frameworks emphasizes how XAI can surpass traditional CNN models in both accuracy and explainability. This is crucial for KC diagnosis, where distinguishing between different types of renal tumors can significantly impact treatment decisions.
Table VII provides a summary of research on ML models enhanced with XAI in KC diagnosis and related analysis. The incorporation of XAI into kidney cancer diagnosis models, as highlighted in the table, significantly enhances the transparency and interpretability of machine learning predictions, which is critical in medical decision-making. Key findings include the successful integration of XAI techniques like SHAP, LRP, and LIME that improve the fairness, understandability, and reliability of model outputs. These XAI-enhanced models facilitate a deeper understanding of how specific features influence diagnostic predictions, thus fostering trust among clinicians and aiding in the effective communication of diagnostic reasoning. For instance, models have achieved high diagnostic accuracies (up to 99.2%) while providing actionable insights into feature relevance, such as the significance of hemoglobin levels in CKD prognosis. This integration addresses a crucial gap in traditional AI approaches by making advanced AI diagnostics more accessible and actionable in clinical settings, particularly enhancing the early detection and treatment planning for kidney-related diseases.
VII-F Adversarial Training
The burgeoning field of adversarial training in DL offers a compelling avenue for enhancing the diagnosis, prediction, and segmentation of KC. This approach leverages the strengths of generative adversarial networks (GANs) and adversarial training methods to address some of the inherent challenges in medical imaging, such as data scarcity, variability in tumor appearance, and the robustness of diagnostic models against adversarial examples. In this respect, Zheng et al. [208] introduce a multi-scale adversarial learning approach for the segmentation of kidney and kidney tumor areas. Their method, which incorporates difficult region-supervised loss and adversarial learning at multiple scales, shows promise in accurately delineating complex tumor shapes and sizes, which are often challenging due to their diverse appearances and fuzzy boundaries. Such an approach is crucial for precise radiotherapy planning and monitoring disease progression, ensuring that treatments are targeted and effective. The MB-FSGAN model discussed by Ruan et al. [209] integrates multi-scale feature extraction and generative adversarial learning to enhance the segmentation and quantification of kidney tumors. This method not only segments but also quantifies tumor indices, providing critical morphometric details that are essential for personalized treatment planning. The adversarial component helps in refining the segmentation accuracy, proving especially useful in clinical settings where precision is paramount. Shan et al. [210] developed SegTGAN, a GAN-based model for autonomous kidney segmentation, which outperformed traditional networks like U-Net and FCN on the Kits19 dataset. The incorporation of multi-scale feature extraction and densely connected blocks in a GAN framework suggests that adversarial training can significantly improve the accuracy of segmentation models, a critical step for effective diagnosis and treatment planning in KC. In the study by Gadermayr et al. [148], the application of adversarial models for image-to-image translation demonstrates a significant advancement in the field of digital pathology segmentation. This approach effectively addresses the variability in staining—a common challenge in pathology—without requiring extensive manual annotations for each characteristic. By enabling stain-independent supervised segmentation and combining it with unsupervised techniques, the study showcases how adversarial training can adaptively manage inherent variabilities in pathological images, leading to improved segmentation outcomes that are crucial for accurate cancer diagnosis and treatment planning. Moving on, Zeng et al. [211] utilize an unsupervised domain adaptation technique to address the segmentation of kidney structures from computed tomography urography images. This method tackles the domain shift problem effectively, demonstrating that adversarial training can bridge the gap between different imaging modalities. By enabling a model to learn deeply from urographic images without direct supervision, this approach provides a way to enhance the 3D reconstruction and visualization of kidneys, which is vital for surgical planning and outcome analysis in KC treatment.
Besides, Uhm et al. [212] propose a unified framework that uses adversarial learning to generate missing CT phases, which are then used to classify KC subtypes. This approach demonstrates how adversarial training can not only enhance data completeness but also improve diagnostic accuracies by enabling the use of multi-phase CT scans that provide comprehensive views of the kidney, essential for accurate diagnosis and staging of cancer. On the other hand, the research by Brinda et al. [213] illustrates the application of Conditional Variational Generative Adversarial Networks (CVGANs) in the diagnosis of chronic kidney disease, which often precedes or accompanies KC. By optimizing the selection of diagnostic features through a metaheuristic algorithm and training the model to classify disease presence more accurately, this study highlights the potential of adversarial networks to refine the prediction capabilities of diagnostic models. The high precision and recall achieved underscore the effectiveness of GANs in enhancing the reliability of computer-aided diagnostic systems. Similarly, the study by Sughasiny et al. [214] introduces the Deep Adversarial Belief Network (DABN), which leverages adversarial training to enhance the classification of renal cell carcinoma (RCC). This approach is significant because RCC is often detected in its late stages, making effective treatment more challenging. The DABN model utilizes adversarial inputs to fine-tune the network’s ability to discern between healthy and cancerous tissue, even when the differences are subtle. This method not only improves accuracy but also addresses the problem of tuning hyperparameters and managing outliers that often affect DL models. By comparing performance with existing methods like E-CNN and FPSO CNN, the DABN demonstrates a higher accuracy, illustrating the potential of adversarial models in refining diagnostic algorithms for early and more precise detection of KC. Moving forward, the work by Liao et al. [215] showcases how DL, specifically through the use of ACWGAN-GP and MobileNetV2, can effectively tackle the problem of imbalanced and insufficient data samples in the classification of chronic kidney disease stages. Adversarial networks, by generating additional training data, can help in creating more balanced datasets that enhance the robustness and generalization ability of diagnostic models. This method could similarly be adapted for KC, where stages or types of cancer could be more accurately classified with augmented training sets.
TABLE VII: Summary of research on ML models enhanced with XAI in KC diagnosis and related analysis.
Ref. ML Model Used Dataset Contribution Description Best Performance Value Limitation
[198] NN-CBR twin system Population health data Implemented a neural network complemented by CBR for explainable predictions of CKD risk, aimed at improving healthcare predictions. Accuracy of 95% Neural networks remain opaque without effective implementation of explainable systems.
[199] VGG16 with XAI (LRP) KUB X-ray images Developed a system for classifying kidney stones using transfer learning and XAI to enhance transparency and decision-making. Accuracy of 97.41% Lacks fairness and explainability without XAI integration.
[200] ConvXGB with SHAP for XAI Survey Lung Cancer dataset Introduced ”DeepXplainer,” a hybrid deep learning model for lung cancer detection and explanation, integrating CNNs and XGBoost with SHAP for explainability. Accuracy of 97.43%, Sensitivity of 98.71%, F1-score of 98.08% The complexity of integrating multiple AI technologies for explainability.
[201] Extreme Gradient Boosting with XAI Clinical features dataset Developed an explainable CKD prediction model using a minimal set of clinical features, optimized for early diagnosis. Accuracy of 99.2% with cross-validation Reliance on a limited number of clinical features which may not capture all nuances of CKD.
[202] DNN with LIME for XAI Healthcare diagnostic data Built a DNN model to predict early-stage CKD with explainability provided by LIME, focusing on interpretability in healthcare applications. Accuracy of 98.75%, ROC_AUC of 98.86% The potential black-box nature of DNN without effective explainability.
[203] Deep learning with GRAD-CAM Robotic nephrectomy videos Utilized deep learning to detect tumors in laparoscopic images, enhancing tumor localization and classification. mAP of 0.974, classification accuracy of 0.84 Specific to images from surgical procedures, may not generalize to other forms of imaging.
[205] Random Forest with Boolean decision trees Data from patients undergoing open PN Predicted AKI post-partial nephrectomy using a Random Forest model optimized for pre-operative parameters. Accuracy of 75% Limited to predictions based on pre-operative data only.
[206] Deep convolutional transfer learning with SHAP Hunan Cancer Hospital and COIL-100 datasets Proposed an RCC prediction model incorporating SHAP for explainability and DCA for evaluating clinical utility. Accuracy of 73.87% on RCC, 99.81% on COIL-100 Lower performance in real-world clinical datasets compared to public datasets.
[207] Explainable AI framework H&E-stained images of ChRCC and RO Developed an explainable framework to differentiate between ChRCC and RO, enhancing diagnostic accuracy and reducing pathologist workload. Diagnostic accuracy of 88.2% The challenge of differentiating similar cancer types based on microscopic features.
[204] CNN with Grad-CAM Abdominal CT images Explored the use of CNNs and visual explanations via Grad-CAM to detect kidney cancer from CT images. Not specified Challenges in handling the diversity of abdominal CT images.
[216] XGBoost with SHAP for XAI Expression data of skin samples Applied XAI to identify significant genes for SCC diagnosis, providing a basis for potential biomarkers. Not specified Specific to non-melanoma skin cancers and dependent on data quality.
[217] Extreme Gradient Boosting with XAI CKD indicators dataset Developed an explainable CKD prediction model, balancing accuracy and explainability using ensemble trees and selected features. Accuracy of 99.2%, cross-validation accuracy of 97.5% Focus on a small set of features, which may not apply universally.
Lastly, in a parallel approach, Hu et al. [218] focus on the issue of image artifacts in prostate cancer diagnosis, which shares similarities with KC imaging challenges. They introduce a Targeted Adversarial Training with Proprietary Adversarial Samples (TPAS) strategy to combat the misleading effects of rectal artifacts in MRI scans. This method specifically generates adversarial samples that mimic these artifacts, training the system to recognize and disregard such misleading information. This adversarial approach enhances the model’s robustness, ensuring that the diagnostic accuracy is not compromised by common imaging anomalies. This strategy underscores the utility of adversarial training in improving the reliability of cancer diagnostics by making the models resilient to specific challenges inherent in medical imaging.
Figure 8: Federated learning for KC.
Table VIII summarizes the relevant research on adversarial training in kidney disease and tumor analysis. It is has been shown that the use of adversarial learning strategies allows for better handling of the inherent variability in tumor appearance, resulting in more precise segmentation outcomes, as demonstrated in studies using models like MSALDS-UNet [208] for fine segmentation tasks. Moreover, adversarial training aids in refining models to detect and segment challenging and less distinct regions, such as fuzzy tumor boundaries, thereby supporting more accurate clinical assessments and potential treatment planning enhancements. These advancements highlight the transformative potential of adversarial training techniques in dealing with the complex imaging characteristics of kidney tumors.
TABLE VIII: Summary of research on adversarial training in KC diagnosis.
Ref. ML Model Used Dataset Contribution Description Best Performance Value Limitation
[215] ACWGAN-GP, MobileNetV2 Renal ultrasound images Developed an auxiliary diagnosis system for CKD stage classification using deep learning. Accuracy up to 90.1% Limited to four stages of CKD.
[208] MSALDS-UNet KiTS21 Proposed a cascade-based approach with multi-scale adversarial learning for kidney and tumor segmentation. Not specified Handling diverse tumor shapes and locations.
[209] MB-FSGAN CT scans of 113 kidney tumor patients Introduced a GAN for simultaneous segmentation and quantification of kidney tumors. Pixel accuracy of 95.7% Manual segmentation variability and time consumption.
[210] SegTGAN Kits19 dataset Developed a GAN-based technique for autonomous kidney segmentation. DSC of 92.28%, ACC of 97.28% Quantitative analysis complexity due to data scarcity.
[212] Fully 3D convolutional networks In-house and external datasets Created a framework for kidney cancer diagnosis with incomplete multi-phase CT. Higher performance than using incomplete data Relies on availability of complete multi-phase CT.
[214] Deep Adversarial Belief Network (DABN) Not specified Utilized Grasshopper Optimization Algorithms for kidney cancer detection. Outperforms existing methods (up to 98.25% accuracy) Hyper-parameter tuning and outlier effects.
[218] Targeted adversarial training with Proprietary Adversarial Samples Not specified Improved prostate cancer classification by mitigating rectal artifact interference. Substantial gains over ordinary training Specific to prostate MRI and adversarial samples.
[148] Image-to-image translation models Patches showing kidney histology Developed unsupervised segmentation approaches for digital pathology. Significant improvements in unsupervised segmentation Requires handling of high image variability.
[211] Unsupervised domain adaptation with 2D networks Clinical urography data Proposed unsupervised kidney segmentation from urographic images. Comparable or better results than supervised methods Adaptation issues across clinical unseen data.
[213] Conditional Variational Generative Adversarial Networks Benchmark CKD dataset Developed a feature selection-based classifier for CKD diagnosis using deep learning. Accuracy of 99.2%, Recall of 98.6% Pre-processing to remove data with absence of values.
VII-G Tranformers

### v2

@@ 6.6 Negative transfer and overfitting
7 Future research directions
Despite achieving high accuracy, prior studies highlight the need for improved determination of KC. Researchers suggest employing less conventional algorithms to enhance early-stage detection. Enhancing TL-based models for robust, unbiased KC identification across stages is crucial. Integrating feature selection with DL models and expanding datasets from various sources may support this goal. Further refinement of algorithms for related classification tasks is also needed. Potential future directions for TL in KC diagnosis are outlined in the following subsections.
7.1 Blockchain technology
In the context of KC diagnosis, TL plays a vital role in enhancing the effectiveness of blockchain technology. TL allows DL models to be trained on large, diverse datasets of medical images, such as CT scans or MRIs, for detecting tumors. Initially, the model is trained on a broad range of cancer data, and later refined using a more specialized dataset focused on KC. This process helps the model to become highly accurate in detecting KC-specific characteristics. Blockchain technology complements TL by providing a secure, transparent, and immutable platform to store and share sensitive medical data. It ensures that patient records are not only protected but also verifiable, offering patients control over their data and granting permissions to authorized medical professionals for analysis. Through blockchain’s decentralized nature, researchers and healthcare institutions can securely share and collaborate on medical data, enabling the use of TL-based models across multiple centers. This integration enhances diagnostic accuracy, reduces bias, and improves patient care outcomes by allowing for more precise, timely, and collaborative diagnoses of kidney cancer [175, 34].
7.2 Adversarial training
The burgeoning field of adversarial training offers a compelling avenue for enhancing the diagnosis, prediction, and segmentation of KC. This approach leverages the strengths of generative adversarial networks (GANs) and adversarial training methods to address some of the inherent challenges in medical imaging, such as data scarcity, variability in tumor appearance, and the robustness of diagnostic models against adversarial examples. Furthermore, TL plays a pivotal role in overcoming the challenge of limited labeled data in medical imaging by enabling the use of pre-trained models on large-scale datasets, which can be fine-tuned for specific tasks like KC detection. TL allows models to capture essential features from general medical imaging datasets and adapt them to the nuances of KC, thus improving diagnostic performance and model generalization. This integration of adversarial training with TL has the potential to significantly enhance the accuracy and robustness of KC diagnosis models, especially in clinical environments with limited annotated datasets [176, 177, 178, 179].
7.3 Compression and denoising based on TL for KC
A future direction for improving KC diagnosis and treatment could involve further advancements in TL. These innovations could enhance both medical image compression and tumor detection processes. Specifically, refining TL techniques for compressing KC images could ensure that critical diagnostic information is preserved while reducing computational and storage demands. Future research could explore more advanced model compression strategies, dynamic loss functions, and real-time data sharing to support collaborative decision-making among healthcare professionals. By combining these strategies, the goal would be to create faster, more accurate, and scalable solutions for detecting and diagnosing KC, leading to better patient care and treatment outcomes [180, 181, 182, 183]. Moreover, employing efficient pretrained denoising DL techniques—such as the one presented in [184]—can significantly improve the quality of medical images prior to analysis, thereby enhancing the accuracy and reliability of KC detection and diagnosis.
7.4 Quantum computing
Quantum computing represents a nascent and burgeoning domain poised to transform our approach to computational challenges, including those within ML, with the potential to revolutionize various aspects of knowledge discovery. Despite being at an early developmental stage, quantum computing harbors the promise of significantly enhancing the efficacy of TL models in the diagnosis of diseases like KC [185]. Employing quantum computing in TL for KC diagnosis entails harnessing quantum algorithms to tackle tasks that would otherwise be impractical with classical computers, notably optimization and linear algebra computations. These algorithms offer avenues to bolster TL model performance by curtailing computational overheads and enhancing model accuracy. Additionally, quantum computing holds promise in mitigating classical computing constraints encountered in medical imaging, such as the demand for substantial computational resources and memory for processing voluminous datasets. By leveraging quantum computing, the duration required for training and inference can be slashed, fostering swifter diagnoses and improved patient outcomes [186, 187, 188].
7.5 DA for clinical data
[…]
7.6 UL, and SSL-based TL for KC
[…]
UL, SSL, and TL for KC diagnosis can provide substantial benefits by enhancing diagnostic accuracy with limited labeled data. unsupervised learning (UL) can uncover hidden patterns and clusters in medical images, identifying previously unnoticed features that distinguish cancerous tissues. SSL can utilize large amounts of unlabeled data by generating pseudo-labels through pretext tasks, reducing the need for manually annotated datasets. TL further amplifies these techniques by enabling models trained on one dataset to be fine-tuned and adapted for KC detection in different clinical settings or patient populations, ensuring better generalization and scalability. Together, these methods can significantly improve early detection, reduce the reliance on expert annotation, and enable the development of more robust, efficient, and scalable AI systems for KC diagnosis [122, 190].
7.7 Federated TL for KC
A promising future research direction involves advancing the integration of TL and federated learning (FL) to tackle the complex issue of identifying kidney abnormalities in medical imaging. This approach could enhance diagnostic capabilities within decentralized healthcare systems while maintaining patient data privacy. Future studies may focus on refining the federated TL architecture to enable collaborative learning across geographically dispersed renal imaging datasets, allowing healthcare providers to share insights without compromising data ownership or confidentiality. Further exploration of this methodology could lead to improved precision in renal abnormality detection by adapting TL models to better capture the unique characteristics of renal imaging data. Additionally, continued research could enhance the efficiency and security of this system, facilitating more effective, privacy-preserving, and decentralized healthcare solutions. By building on this convergence of deep learning, federated learning, and TL, future work could significantly transform the landscape of medical image analysis, fostering a new era of collaborative, secure, and adaptive healthcare [191, 192].
7.8 Self-paced learning for KC
TL has proven to be a powerful approach for improving medical diagnostic systems, especially in the complex field of KC diagnosis. TL leverages knowledge gained from pre-trained models on large datasets to enhance the performance of models on specific tasks, such as KC detection. By incorporating a strategy that focuses on progressively more complex or informative cases as the model improves, TL adapts to the nuances of the data, helping to address challenges like data variability and imbalance. This adaptive learning process can greatly benefit KC diagnosis, allowing the model to effectively handle the complexities of medical imaging and improve diagnostic accuracy over time [193].
7.9 Tranformers and Large language models for KC